OpenStudy (anonymous):

How to find the square root of 39

OpenStudy (anonymous):

you have to use a calculator. the answer would be 6.2

OpenStudy (anonymous):

I know what the answer is but i want to know how to find it?

OpenStudy (anonymous):

Please help its important

OpenStudy (anonymous):

You can use Newton's method.

OpenStudy (amoodarya):

\[\sqrt{x^{2}+a }\approx x+\frac{ a }{ 2x }\\ \sqrt{39}=\sqrt{36+3}=\sqrt{6^2+3}\\x=6\\a=3\\sqrt{6^{2}+3 }\approx 6+\frac{ 3 }{ 2(3)}\]

OpenStudy (anonymous):

I am in 9th grade

Parth (parthkohli):

lol

OpenStudy (anonymous):

Okay, so first you realize: \[ 36<39<49 \\ 6<\sqrt{39}<7 \]

OpenStudy (anonymous):

Next you can try guessing and checking: \[ 6.5^2 = 42.25 \]Too high... \[ 6.2^2 = 38.44 \]Too low... \[ 6.25^2 = 3.90625 \]Getting closer

OpenStudy (anonymous):

I didn't get u

OpenStudy (anonymous):

that's cause no teacher would expect u to get that without a calculator...

OpenStudy (anonymous):

see as i said before i want to know the method not the answer

OpenStudy (anonymous):

those are the only ways possible.. what class is this for?

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

Heron's Formula

OpenStudy (phi):

There is a method to find the square root "by hand", but I don't think it is taught nowadays. either leave the answer written as \( \sqrt{39}\) or use a calculator

OpenStudy (phi):

FYI, here is an example of the algorithm (notice it is a bit complicated) http://johnkerl.org/doc/square-root.html

OpenStudy (anonymous):

Long and tedious method. Anyway, here it is.